OpenStudy (cherrilyn):

use the comparison test to determine whether or not the integral converges...

OpenStudy (cherrilyn):

\[\int\limits\limits_{1}^{\infty} 1/x ^{4}+e ^{x} dx\]

OpenStudy (anonymous):

or you can use the integral test in this case

OpenStudy (anonymous):

\[\lim_{R \rightarrow \infty} \int\limits_{1}^{R} (\frac{1}{x^4} + e^x )dx\]

OpenStudy (anonymous):

and integrate normally, if you end up witha finite number = converges, but if you end up with infinity = diverges ^_^

OpenStudy (anonymous):

give it a try now :)